A bobsled leaves a horizontal track with a speed of 30m/s. it then moves onto a horizontal surface lightly sprinkled with sawdust and comes to a stop after 150m. find the coefficient of friction
1
Expert's answer
2018-04-26T10:20:08-0400
From the conservation of energy:
(mv^2)/2=μmgd
μ=v^2/2gd
The coefficient of friction:
μ=〖30〗^2/(2(10)(150))=0.3.